Dan Dayle started a business by issuing an $92,000 face value note to First State Bank on January 1, 2018. The note had an 7 percent annual rate of interest and a five-year term. Payments of $22,438 are to be made each December 31 for five years.

Required

  1. What portion of the December 31, 2018, payment is applied to interest expense and principal?

  2. What is the principal balance on January 1, 2019?

  3. What portion of the December 31, 2019, payment is applied to interest expense and principal? (Round your answers to the nearest dollar amount.)

a. Interest expense
Principal
b. Principal balance
c. Interest expense
Principal
Answer #1

a) Interest expense = 92000*7% = 6440

Principal = 22438-6440 = 15998

b) Principal balance = 92000-15998 = 76002

c) Interest expense = 76002*7% = 5320

Principal = 22438-5320 = 17118